OpenStudy (anonymous):

what is the vaule of a1 for a geometric sequence with a4 =40 and a6 =160?

OpenStudy (anonymous):

Find a5 first by taking Geometric Mean: \[\large a_5 = \sqrt{a_4 \times a_6} = a_5 = \sqrt{40 \times 160} = a_5 = 80\]

OpenStudy (anonymous):

Now find r: \[r = \frac{a_5}{a_4} = \frac{80}{40} = 2\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

So: Third term will be: 40/2 = 20 Second term will be : 20/2 = 10 And first term will be : 10/2 = 5. \(\color{green}{\textbf{First Term is : 5..}}\)